# Nonlocal mass creation by metricized accelerations in Euclidean matterspace

### v1

@@ 2 Accelerations define field masses and monistic action
The tensor curvature $B_{\mu\nu}\equiv\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu}$ $=R_{\mu\nu}+\theta_{\mu\nu}$, where $R_{\mu\nu}\equiv\partial_{\lambda}\Gamma^{\lambda}_{\mu\nu}-\partial_{\nu}\Gamma^{\lambda}_{\mu\lambda}+\Gamma^{\lambda}_{\mu\nu}\Gamma^{\rho}_{\rho\lambda}-\Gamma^{\lambda}_{\mu\rho}\Gamma^{\rho}_{\nu\lambda}$, $g^{\mu\nu}\theta_{\mu\nu}\equiv 0$, $g^{\mu\nu}B_{\mu\nu}=g^{\mu\nu}R_{\mu\nu}$ and $B_{\mu\nu}\neq R_{\mu\nu}$, is based in the novel geometric analyses on the covariant derivatives of four-acceleration $a_{\mu}$ under strict 3D-flattness, $g_{i}g_{j}-g_{ij}\equiv\delta_{ij},g_{\mu}\equiv g_{o\mu}/g_{o}$, $g^{ij}\equiv-\delta^{ij}$, and adaptive changes of local time, $d\tau\equiv g_{\mu}dx^{\mu}/c\neq dx^{o}/c$. For static self-organization with spherical symmetry, where $u_{i}\equiv\gamma(g_{i}-\beta_{i})=0,u^{i}\equiv\beta^{i}\gamma=0$, $g_{oi}=0$, $g^{oi}=0$, $\gamma\equiv 1/{\sqrt{1-\beta^{2}}}=1$, one can use the already derived [12] solution $g_{o}^{2}u^{o}=u_{o}\equiv g_{o}\gamma=g_{o}=r/(r+r_{o}),$ which results in $a^{o}\equiv u_{\nu}f^{\nu o}=0$, $a_{o}\equiv u^{\nu}f_{\nu o}=0$, $a^{i}\equiv u_{\nu}f^{\nu i}=u_{o}f^{oi}=u_{o}g^{o\lambda}g^{i\rho}f_{\lambda\rho}=u_{o}g^{oo}g^{ij}f_{oj}=\delta^{ij}\partial_{j}lng_{o}=\delta^{ij}\Gamma_{oj}^{o}=r_{o}x^{i}/[r^{2}(r+r_{o})]$ and $a_{i}\equiv u^{\nu}f_{\nu i}=u^{o}f_{oi}=-\partial_{i}lnu_{o}=-\Gamma^{o}_{oi}=-\delta_{ij}a^{j}$. Thus, the 2007 solution [12] for the strong-field interactions with probe (external) masses, $w_{i}(r)\equiv-Gmx_{i}/[r^{2}(r+r_{o})]\equiv a_{i}(r)c^{2}$, is determined by the Einsteinian 3-acceleration $a_{i}(r)$ of the static material densities within their correlated distribution in the volumetric mass integral.
@@
The relativistic four-acceleration $a_{\mu}\equiv g_{\mu\nu}a^{\nu}$ is a primary characteristic of the local creation and dynamic self-distribution of massive densities in the nonlocal mass-energy $mc^{2}=const$. The action (2) with $\nabla_{\mu}a^{\mu}(x)$ would be dropped in the dualistic mechanics for localized masses near the coordinate origin due to the Gauss theorem for the distant hypersurface. In this traditional way, the mass-generation mechanism associated with $\nabla_{\mu}a^{\mu}(x)>0$ in the monistic mechanics of massive fields was dropped due to the assumed dualism of matter with massless geometric fields referenced by Newtonian empty space.
@@
In the monistic mechanics for quasi-isolated (geometrized) distributions both geometric fields and inertial currents are defined by the local self-acceleration $a^{\mu}$ of matterspace with adaptive local times. The 3-acceleration $a^{i}\equiv g^{i\nu}a_{\nu}\equiv u_{\nu}f^{\nu i}\neq 0$ is responsible for everything in static self-organizations ($g^{oi}=0,g^{oo}=1/g_{oo}>1,a^{o}=0$, $a^{i}=-\delta^{ij}a_{j}$), including the generation of both active, $\mu_{a}=c^{2}\partial_{i}a^{i}/4\pi G$, and passive, $\mu_{p}=c^{2}a^{i}(\partial_{i}lng_{o})/4\pi G$, mass densities or Maxwell-type currents of mechanical energy. The spatial 3-divergence is well known as a local density of (active) charges in Coulomb forces.
@@
In mechanics, the acceleration three-vector $a_{i}=-\delta_{ij}a^{j}$ works as the interaction field for the probe mass. The temporal divergence $\nabla_{o}a^{o}=\Gamma_{oi}^{o}a^{i}=-a_{i}a^{i}>0$ seems to be a local self-action associated with the metric time auto-changes $cd\tau/dx^{o}={\sqrt{g_{oo}}}=r/(r+r_{o})\neq const$. The Minkowski metric with $g_{oo}=1,g_{ij}=-\delta_{ij}$ has no local self-accelerations, $a_{o}=a^{o}=a_{i}=a^{i}=0$, $\nabla_{o}a^{o}=\nabla_{i}a^{i}=0$, and can produce neither the passive mass density $\mu_{p}$, nor the active $\mu_{a}$. Following the original initiative of Grossmann in 1912, the co-authors made only temporal changes in the Minkowski 4-interval, $ds=\sqrt{c^{2}d\tau^{2}(g_{\mu\nu})-\delta_{ij}dx^{i}dx^{j}}$, preserving Euclidean 3-space [14]. Metric time dilation can quantitatively explain all post-Newtonian corrections (light bending, Shapiro echo delay, perihelion precession, etc.) and to propose the acceleration mechanism for mass generation. It now appears that Grossmann reasonably resisted signing Einstein’s later manuscripts with the curvature of 3-space.
@@
3 Non-local auto-dynamics of flat matterspace with adaptive time rate
@@
The time-varying self-organization of correlated densities in nonlocal matterspace maintains everywhere $u_{\mu}u^{\mu}\equiv 1$ for the 4D geometry of 3+1 manifolds with adaptive time rates in the spatial continuum. Such self-consistent geometry never distorts the Euclidean 3D subgeometry in adaptive self-movements of the nonlocal unity. The theoretical requirement to preserve Euclidean 3-space by six inherent symmetries $g^{ij}\equiv-\delta^{ij}$ or $g_{ij}\equiv g_{i}g_{j}-\delta_{ij}$ reduces ten Hilbert variables $g^{\mu\nu}=g^{\nu\mu}$ [11] to four, $g^{\mu o}=g^{o\mu}$. One can eliminate the spatial variations in the Hilbert formalism, $T_{ij}\delta g^{ij}\equiv T_{ij}\delta(-\delta^{ij})=0$ in spite of $T_{ij}(x)\neq 0$.
@@
The massive field self-organization with $T^{\nu}_{o}(x^{i},x^{o})=T_{\nu}^{o}(x^{i},x^{o})=0$ and $T^{i}_{j}(x^{i},x^{o})\neq 0$ means that there is no 4D energy continuum for observations. Only 3D space $x^{i}$ is continuously filled by the measurable mass-energy, while $x^{o}$ is not the continuum coordinate. Rather, it is a world running parameter to form a specific metric time for each quasi-isolated system within the common spatial superposition under universal 3D geometry. If the external forces are neglected, the quasi-closed system is described by metric-kinetic flows of mass-energy in flat space. The corresponding components of the symmetric metric tensor $g_{\mu\nu}$ and the affine connections $\Gamma^{\lambda}_{\mu\nu}$ can be expressed in terms of the metric 4-potential $\varphi_{\!{}_{{}_{G}}}g_{\mu}$ with $g_{\mu}\equiv g_{o\mu}/{\sqrt{g_{oo}}}$:
@@
Not arbitrary, but physically admissible transformations between 3+1 manifolds with inherent metric symmetries for Euclidean subgeometry, where $|g_{ij}(x)|$ $=1$ and ${\sqrt{{-g}(x)}}=g_{o}(x)$, are responsible for the tensor covariance in equations (4). The adaptive dilation $cd\tau(x)/dx^{o}\equiv g_{\mu}dx^{\mu}/dx^{o}$ of the physical time rate $d\tau$ in the geometric organization is not a primary entity to define the material processes, but a secondary function or observable phenomenon in the world picture and its description.
@@
The covariant divergence of $T^{\nu}_{\mu}(x)$ should vanish in the Hilbert formalism due to the Noether theorem. The geometrized dynamics of the correlated continuum of inertial fields obeys the contracted Bianchi identity with the curvature tensor ${B}_{\mu\nu}(x)\equiv\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu}$ employed in terms of the relativistic four-accelerations $a_{\mu}(x)\equiv g_{\mu\nu}(x)a^{\nu}(x)$:
$\cases{\nabla_{\nu}T^{\nu}_{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}\nabla_{\nu}(\delta_{\mu}^{\nu}\frac{B}{2}-B_{\mu}^{\nu})\equiv\frac{c^{4}}{8\pi G}(\nabla_{\mu}\nabla_{\nu}a^{\nu}-\nabla_{\nu}\nabla_{\mu}a^{\nu}-\nabla_{\nu}\nabla^{\nu}a_{\mu})\cr\equiv-\frac{c^{4}}{8\pi G}(a^{\nu}R_{\nu\mu}+\nabla_{\nu}\nabla^{\nu}a_{\mu})=0,\cr a^{\nu}(\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu})+\nabla_{\nu}\nabla^{\nu}a_{\mu}=0,\ $if$\ a^{\nu}R_{\nu\mu}=a^{\nu}B_{\nu\mu},\ $and$\ \cr\nabla^{\nu}\nabla_{\nu}a_{o}=-2a^{o}\nabla_{o}a_{o},\ \nabla^{\nu}\nabla_{\nu}a_{i}=-a^{j}(\nabla_{i}a_{j}+\nabla_{j}a_{i})\ $for$\ B_{oj}=0.}$ (7)
One can rewrite the Euclidean space 3-balances in the 4-vector equation (6) in another equivalent form by using the geometric constraints ${\sqrt{-g(x)}}=g_{o}(x),g_{ij}(x)=g_{i}(x)g_{j}(x)-\delta_{ij}$ and the variational consequence $T^{\mu}_{o}(x)=T^{\mu o}(x)$ $=0$ for the correlated dynamics,
$\nabla_{\nu}T^{\nu}_{i}=\frac{\partial_{\nu}({\sqrt{-g}}T^{\nu}_{i})}{\sqrt{-g}}-\frac{T^{\rho\lambda}}{2}\partial_{i}g_{\rho\lambda}=\partial_{j}T^{j}_{i}+T^{j}_{i}\partial_{j}lng_{o}-\frac{T^{kj}}{2}\partial_{i}(g_{k}g_{j})=0.$ (8)
@@
Inhomogeneous mass distributions cannot exist in static states in Newtonian mechanics and gravitation. These dualistic theories were developed for energy and momentum exchange between separated bodies in the void, not for nonlocal correlations of the elastic whole as a wave function continuum in quantum mechanics. The radial mass-energy densities of the static matterspace [12],
$\mu_{p}(r)c^{2}=\mu_{a}(r)c^{2}=\frac{c^{4}B(r)}{16\pi G}=\frac{Gm^{2}c^{4}}{4\pi r^{2}(c^{2}r+Gm)^{2}}\Rightarrow\frac{c^{4}}{4\pi Gr^{2}},$ (9)
has the universal curvature near the origin, $lim_{r\rightarrow 0}B(r\ll Gm/c^{2})$ $={4}/{r^{2}}$. For relatively large distances from the center of spherical symmetry, $r\gg Gm/c^{2}$, the centripetal gradients $c^{4}\partial_{i}B(r)/16\pi G<0$, are proportional to the mass integral $m$ and would predict the gravitational collapse for interacting partners or probe bodies.
@@
In the monistic geometrodynamics of the nonlocal whole, there are no partners. Instead, there is a self-action of correlated densities. The corresponding self-accelerations, or counterforces $c^{4}\nabla_{\nu}B^{\nu}_{i}/8\pi G\equiv c^{4}(\nabla_{\nu}\nabla_{i}a^{\nu}+\nabla_{\nu}\nabla^{\nu}a_{i})/8\pi G$ after their substitution on the right hand side in (6), prevent the centripetal collapse of inhomogeneous energy profiles with $(\partial_{i}B)c^{4}/16\pi G\neq 0$. The covariant counterbalance for non-equilibrium oscillations around equilibrium profiles and for stationary organisations can be easily verified in (6)-(7) under the static self-organization of relativistic 4-velocities $cu_{\mu}\equiv g_{\mu\nu}cu^{\nu}\equiv\{cg_{o}\gamma;-c\gamma(\beta_{i}-g_{i})/g_{o}\}$ =$=\{c{\sqrt{g_{oo}}};0\}$ and 4-accelerations $a_{\mu}\equiv g_{\mu\nu}a^{\nu}\equiv c^{2}u^{\nu}f_{\nu\mu}=\{0;-c^{2}x_{i}r_{o}/[r^{2}(r+r_{o})]\}$, $a^{\mu}=\{0;+c^{2}x^{i}r_{o}/[r^{2}(r+r_{o})]\}$ of the generated mass-energy densities (8) with their metric consequences $g_{oo}=1/g_{oo}=r^{2}/(r+r_{o})^{2},g_{oi}=g^{oi}=0,g_{ij}=-\delta_{ij},g^{ij}=-\delta^{ij}$.
4 Shannon information in the yin-yang balance of distributed energies

### v2

@@ 2.3 Inherent space-time symmetris for Euclidean 3-geometry
For static self-organisation of massive fields with spherical symmetry, where $\gamma\equiv(1-\beta^{2})^{-1/2}=1,u_{o}={\sqrt{g_{oo}(r)}}\equiv g_{o}(r)=r/(r+r_{o}),g_{i}\equiv g_{oi}/g_{o}=0,g_{oi}=0,g^{oi}=0,g^{ij}=-\delta^{ij}$, $g_{ij}=-\delta_{ij}$, $u_{i}=u^{i}=0$, one can use the already known metric solutions: $f_{oi}=-\partial_{i}g_{o}=-(x_{i}r_{o}/r)/(r+r_{o})^{2},f^{oi}=g^{oo}g^{ij}f_{oj}=\delta^{ij}x_{i}r_{o}/r^{3},(-f_{oi}f^{oi})=r^{2}_{o}/[r(r+r_{o})]^{2}=g_{o}i^{o},i^{o}=r_{o}^{2}/[r^{3}(r+r_{o})]$. These metric fields with radial inertial densities correspond to $a^{o}\equiv u_{i}f^{oi}=0,a_{o}\equiv u^{i}f_{io}=0$, $a^{i}=u_{o}f^{oi}\equiv u_{o}g^{o\lambda}g^{\rho i}f_{\lambda\rho}=u_{o}g^{oo}g^{i\rho}f_{o\rho}=-g_{o}g^{oo}f_{oj}\delta^{ij}=\delta^{ij}\partial_{j}lng_{o}=\delta^{ij}\Gamma^{o}_{oj}=x_{i}r_{o}/[r^{2}(r+r_{o})]$, and $a_{i}=u^{o}f_{oi}=-\partial_{i}lnu_{o}=-\Gamma^{o}_{oi}=-\delta_{ij}a^{j}$. Thus, the 2007 solution [16] for the strong-field interactions with probe (external) masses, $w_{i}(r)\equiv-Gm\delta_{ij}x^{j}/[r^{2}(r+r_{o})]=a_{i}(r)c^{2}$, is determined by the Einsteinian 3-acceleration $c^{2}a_{i}(r)$ of the static material densities within their correlated distribution in the volumetric mass integral $m$.
2.4 Primacy of local accelerations for field matter observations by probe bodies
@@
The relativistic four-acceleration $a_{\mu}\equiv g_{\mu\nu}a^{\nu}$ is a primary characteristic of the local creation and dynamic self-distribution of massive densities in the nonlocal mass-energy $mc^{2}=const$. The action (2) with $\nabla_{\mu}a^{\mu}(x)$ would be dropped in the dualistic mechanics for localized masses near the coordinate origin due to the Gauss theorem for the distant hypersurface. The nonlocal mass generation mechanism associated with $\nabla_{\mu}a^{\mu}(x)>0$ in the monistic action of massive fields has no analogues in dualastic Lagrangians of localized masses and massless fields on distant surfaces.
@@
In the monistic mechanics for quasi-isolated (geometrized) distributions both geometric fields and inertial currents are defined by the local self-acceleration $a^{\mu}$ of matterspace with adaptive local times. The 3-acceleration $a^{i}\equiv g^{i\nu}a_{\nu}\equiv u_{\nu}f^{\nu i}\neq 0$ is responsible for everything in static self-organizations ($g^{oi}=0,g^{oo}=1/g_{oo}>1,a^{o}=0$, $a^{i}=-\delta^{ij}a_{j}$), including the generation of both active, $\mu_{a}=c^{2}\partial_{i}a^{i}/4\pi G$, and passive, $\mu_{p}=c^{2}a^{i}(\partial_{i}lng_{o})/4\pi G$, mass densities or Maxwell-type currents of mechanical energy.The spatial 3-divergence is well known as a local density of (active) charges in Coulomb forces. The absence of matterspace accelerations relative to the local observer implies that neither the passive mass density nor the active mass density cannot emerge in the geodesically falling elevator for the practical detection of dense materspace for non-geodesically moving observers.
@@
In mechanics, the acceleration three-vector $a_{i}=-\delta_{ij}a^{j}$ works as the interaction field for the probe mass. The temporal divergence $\nabla_{o}a^{o}=\Gamma_{oi}^{o}a^{i}=-a_{i}a^{i}>0$ seems to be a local self-action associated with the metric time auto-changes $cd\tau/dx^{o}={\sqrt{g_{oo}}}=r/(r+r_{o})\neq const$. The Minkowski metric with $g_{oo}=1,g_{ij}=-\delta_{ij}$ has no local self-accelerations, $a_{o}=a^{o}=a_{i}=a^{i}=0$, $\nabla_{o}a^{o}=\nabla_{i}a^{i}=0$, and can produce neither the passive mass density $\mu_{p}$, nor the active $\mu_{a}$. Following the original initiative of Grossmann in 1912, the co-authors made only temporal changes in the Minkowski 4-interval, $ds=\sqrt{c^{2}d\tau^{2}(g_{\mu})-\delta_{ij}dx^{i}dx^{j}}$, preserving Euclidean 3-space [18]. Again, metric time dilation can quantitatively explain all post-Newtonian corrections (light bending, Shapiro echo delay, perihelion precession, etc.) and to propose the acceleration mechanism for mass generation. It now appears that Grossmann was reasonably reluctant to endorse or sign Einstein’s later manuscripts with the hypothetical curvature of 3-space in the advised 4D pseudo-Riemann geometry.
@@
3 Non-local geometrodynamics of Euclidean matterspace with adaptive time rate
3.1 Equivalence of mass densities from four variational equations
In a flat matterspace characterized by affine connections over the whole volume, the nonlocal self-organization of correlated densities in the closed system always satisfies the basic conditions: $u_{\mu}u^{\mu}\equiv 1$ and $u_{\mu}a^{\mu}\equiv 0$. The specific 4D geometry of each closed (elementary) system preserves local properties without distorting the universal 3D Euclidean geometry for the hierarchy of superimposed 3-volumes of different geometric systems. The requirement to preserve the universal subgeometry of 3D physical spaces for the multi-system universe with nonlocal correlations leads to the reduction of the ten variables in the Hilbert formalism to four. The effective geometrodynamics of each metric system, in particular with respect to its mass-energy conservation, is based on the correlation of the specific metric potentials $g_{\mu}c^{2}$ and the specific adaptivity of the local time rates $cd\tau/dx^{o}\neq 1$.
@@ 3.1 Equivalence of mass densities from four variational equations
Again, time-varying and equilibrium self-organizations of correlated densities in nonlocal matterspace maintains everywhere the curved 4D geometry of 3+1 manifolds with adaptive time rates for the flat spatial continuum. Such self-consistent geometry never distorts the Euclidean 3D subgeometry in adaptive self-movements within the nonlocal unity. The theoretical requirement to preserve Euclidean 3-space by six inherent symmetries $g^{ij}\equiv-\delta^{ij}$ or $g_{ij}\equiv g_{i}g_{j}-\delta_{ij}$ reduces ten Hilbert variables $g^{\mu\nu}=g^{\nu\mu}$ [11] to four, $g^{\mu o}=g^{o\mu}$. One can eliminate the spatial variations in the Hilbert formalism, $T_{ij}\delta g^{ij}\equiv T_{ij}\delta(-\delta^{ij})=0$ in spite of $T_{ij}(x)\neq 0$.
@@
The trace of the energy-momentum tensor in the monistic field theory is equal not to zero, but to the positive sum of two equal densities: $\mu_{a}(x)c^{2}+\mu_{p}(x)c^{2}=T^{\mu}_{\mu}(x)>0.$ The time-varying mass equality, $\mu_{a}(x^{i},x^{o})=\mu_{p}(x^{i},x^{o})$ for the active density $\mu_{a}(x)\equiv$ ${\varphi_{\!{}_{{}_{G}}}^{2}}\nabla_{i}a^{i}(x)/{4\pi}$ $\equiv c^{2}B^{i}_{i}(x)/8\pi G$ and the passive one $\mu_{p}(x)\equiv{\varphi_{\!{}_{{}_{G}}}^{2}}\nabla_{o}a^{o}(x)/{4\pi}$ $\equiv c^{2}B^{o}_{o}(x)/8\pi G$, was derived in (4) from the Hilbert variations as the Lagrange equation of motion, but not postulated from the General Theory of Relativity as a principle. The positive scalar densities continuously fill the correlated 3-volume with the fixed mass-energy integral, $const=$ $\int d^{3}x{\sqrt{-g}}[\mu_{a}(x^{i},x^{o})+\mu_{p}(x^{i},x^{o})]$.
The spatial density $B(x^{i},x^{o})>0$ is the scalar curvature for the 4D self-organization. This density provides the positive mass-energy integral to the observable 3D world with the universal 3D subgeometry. The framework presented reflects an inextricable link between active and reactive mass densities, based on geometric principles that maintain the integrity of physical laws across different frames of reference. The quantitative equality of active and reactive mass densities, which are different functions of the relativistic acceleration, underlines the essential nature of metric transformations and the correlated geometrodynamics. Monistic matterspace with universal 3D geometry and adaptive time unifies field masses and thier energy densities in a nonlocal material volume of the pseudo-Riemannian 4D organization.
3.2 Zero energy integral for a closed mechanical self-assembly
The adaptive 4D manifolds with specific times in different auto-organizations of quasi-isolated systems maintain zero variational densities $T_{\mu o}(x)=0$ resulting in $T^{\nu}_{o}(x)=g^{\nu\mu}(x)T_{\mu o}(x)=0$, while $T^{\mu o}(x)\equiv g^{o\nu}T^{\mu}_{\nu}=g^{oi}T_{i}^{\mu}\neq 0$. The system mechanical energy for nonlocally correlated material densities vanishes after 3D integration over the hypersurfaces,
@@ 3.2 Zero energy integral for a closed mechanical self-assembly
This finding from the Hilbert variations with Euclidean 3D section seems strange for inertial fields with only positive mass densities. However, the zero energy integral of a closed mechanical system with self-gravity was intuitively discussed (by P. Dirac, P. Jordan, R.J. Glauber and others) in the 1930s-1950s period. Dualistic physics exploited the idea of negative binding energy in isolated systems like a secluded star. But General Relativity with warped 3-space failed in quantitative calculations of strong field energy balances for the sum of interacting masses. The flat-space result (5) corresponding to the equivalence of active and passive mass densities, is deciphered in the last section by analytical examples with the Shannon information within a continuous energy distribution.
The massive field self-organization with $T^{\nu}_{o}(x^{i},x^{o})=0$ and $T_{i}^{\nu}(x^{i},x^{o})\neq 0$ means that there is no 4D energy continuum for observations. Only 3D space $x^{i}$ is continuously filled by the measurable mass-energy, while $x^{o}$ is not the continuum coordinate. Rather, it is a world running parameter to form a specific metric time for each quasi-isolated system within the common spatial superposition under universal 3D geometry. If the external forces are neglected, the quasi-closed system is described by metric-kinetic flows of mass-energy in flat space. Both the corresponding components of the symmetric metric tensor $g_{\mu\nu}$ and the affine connections $\Gamma^{\lambda}_{\mu\nu}$ can be expressed in terms of $g_{\mu}\equiv g_{o\mu}/{\sqrt{g_{oo}}}$ or the metric 4-potential $\varphi_{\!{}_{{}_{G}}}g_{\mu}$ when 3D continuum of field masses is flat:
@@
These expressions form the basis for determining how the metric behaves under transformations that satisfy the requirements of tensor covariance. Not arbitrary, but physically admissible transformations between 3+1 manifolds with inherent metric symmetries for Euclidean subgeometry, where $|g_{ij}(x)|$ $=1$ and ${\sqrt{{-g}(x)}}=g_{o}(x)$, are responsible for the covariance in equations (4).The adaptive dilation $cd\tau(x)/dx^{o}\neq 1$ of the proper (physical) time rate in the geometric organization is not a primary entity to define the material processes, but a secondary function or observable phenomenon in the world picture and its description. Again, the physical time does not act as a primary driver of material processes. Rather, it emerges from the metric organization of the specific 4D manifold, acting as a secondary consequence of its geometrodynamics.
3.3 Geometrodynamic counterbalance in non-equilibrium self-organisations
@@
The covariant divergence of $T^{\nu}_{\mu}(x)$ should vanish in the Hilbert formalism due to the Noether theorem. The geometrized dynamics of the correlated continuum of inertial fields obeys the contracted Bianchi identity with the curvature tensor defined as ${B}_{\mu\nu}(x)\equiv\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu}$, $B_{\mu}^{\nu}\equiv g^{\nu\lambda}B_{\mu\lambda}$, and $B\equiv 2\nabla_{\mu}a^{\nu}$:
$\nabla_{\nu}T^{\nu}_{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}\nabla_{\nu}(\delta_{\mu}^{\nu}\frac{B}{2}-B_{\mu}^{\nu})\equiv\frac{c^{4}}{8\pi G}(\nabla_{\mu}\nabla_{\nu}a^{\nu}-\nabla_{\nu}\nabla_{\mu}a^{\nu}-\nabla_{\nu}\nabla^{\nu}a_{\mu})=0.$ (8)
This covariant consequence of correlated stresses in any pseudo-Riemannian space-time reads for local 4-accelerations $a_{\mu}(x)$ of emerging mass densities as the geometric feedback $\nabla_{\nu}\nabla^{\nu}a_{\mu}=-R_{\mu\nu}a^{\nu}$.
One can study geometrodynamic balances along 3 spatial axes $x^{i}$ in the material continuum by rewriting the tensor contraction (7) in another equivalent form using 3D geometric constraints ${\sqrt{-g(x)}}=g_{o}(x)$ and the variational consequence $T^{\nu}_{o}(x)$ $=0$ for non-local self-assembly, where $T^{\nu}_{i}(x)\equiv g^{\nu\mu}T_{i\mu}(x)=g^{\nu j}T_{ij}(x)$ $\neq 1$ and $g_{o}T^{oo}\equiv g_{o}g^{\nu o}T_{\nu}^{o}(x)=g^{j}T^{o}_{j}(x)$ $\neq 1$:
$\nabla_{\nu}T^{\nu}_{i}=\frac{\partial_{\nu}({\sqrt{-g}}T^{\nu}_{i})}{\sqrt{-g}}-\frac{T^{\mu\nu}}{2}\partial_{i}g_{\mu\nu}=\frac{\partial_{\nu}({g_{o}}g^{\nu j}T_{ij})}{g_{o}}$ (9)
$-\delta^{kj}g_{k}T^{o}_{j}\partial_{i}g_{o}-T^{oj}\partial_{i}(g_{o}g_{j})-\frac{T^{kj}}{2}\partial_{i}(g_{k}g_{j})=0.$ (10)
The interplay of the derivatives with four metric potentials suggests how the geometry is embedded in the adaptive geometrodynamics of the emerging mass densities in the energy-momentum tensor. The covariant counterbalance in (7) or (8) for non-equilibrium oscillations of elementary matterspace around the equilibrium profile of inertial densities implies that the geometrodynamic self-assembly maintains stability under perturbations. The tensor expressions for the 4-velocities and 4-accelerations of the generated mass densities show how these coordinate fields are intertwined with the emerging 4D geometry of a specific hierarchical system.
@@
Inhomogeneous mass distributions cannot exist in static states in Newtonian dual mechanics and gravitation. This dualistic theory was developed for energy and momentum exchanges between separated partners in a void, rather than for nonlocally controlled self-accelerations of the elastic whole as a nonlocal continuum in quantum mechanics. The radial mass-energy densities of the static matterspace, as described in reference [16], are given by:
$\mu_{p}(r)c^{2}=\mu_{a}(r)c^{2}=\frac{c^{4}B(r)}{16\pi G}=\frac{Gm^{2}c^{4}}{4\pi r^{2}(c^{2}r+Gm)^{2}}\Rightarrow\frac{c^{4}}{4\pi Gr^{2}}.$ (11)
This continual solution shows that there is universal curvature near the coordinate origin, $lim_{r\rightarrow 0}R(r\ll Gm/c^{2})$ $={4}/{r^{2}}$. For relatively large distances from the center of spherical symmetry, $r\gg Gm/c^{2}$, the radial mass densities are proportional to $m^{2}/r^{4}$, which is the self-action of Newtonian interaction fields.
@@
In the monistic physics of the nonlocal whole, there are no partners; instead, there is a self-action of correlated fields. The non-Newtonian self-tensions in (7) along spatial directions, $-c^{4}\nabla_{\nu}B^{\nu}_{i}/8\pi G\equiv-c^{4}(\nabla_{\nu}\nabla_{i}a^{\nu}+\nabla_{\nu}\nabla^{\nu}a_{i})/8\pi G$, prevent the centripetal collapse of inhomogeneous energy profiles with $(\partial_{i}B)c^{4}/16\pi G$ $\neq 0$. The covariant counterbalance for non-equilibrium oscillations around equilibrium profiles and for stationary organisations can be easily verified in (7)-(8) under the static self-organization of relativistic 4-velocities $cu_{\mu}\equiv g_{\mu\nu}cu^{\nu}\equiv$$\left\{cg_{o}\gamma;-c\gamma\left(\beta_{i}-g_{i}\right)/g_{o}\right\}\Rightarrow\left\{c\sqrt{g_{oo}};0\right\}$ and GR 4-accelerations $c^{2}a_{\mu}\equiv c^{2}g_{\mu\nu}a^{\nu}\equiv c^{2}u^{\nu}f_{\nu\mu}$$\Rightarrow\left\{0;-c^{2}x_{i}r_{o}/\left[r^{2}\left(r+r_{o}\right)\right]\right\},a^{\mu}=\left\{0;+c^{2}x^{i}r_{o}/\left[r^{2}\left(r+r_{o}\right)\right]\right\}$ of the generated mass-energy densities in (9) with the metric consequences $g_{oo}=1/g_{oo}=r^{2}/(r+$ $\left.r_{o}\right)^{2},g_{oi}=g^{oi}=0,g_{ij}=-\delta_{ij},g^{ij}=-\delta^{ij}$.
The monistic framework presented elucidates how adaptive time rates and nonlocal correlations can explain phenomena traditionally handled within a Newtonian dualism, while simultaneously preserving the integrity of the spatial geometry. The new theoretical construct opens avenues for understanding cosmic dynamics without relying on the conventional partnership of interacting bodies in dualistic frameworks. The conceptual monism of continuous material organizations can potentially guide future research in geometrodynamics for the nonlocal organization of self-gravitating systems.
4 Shannon information in the Yin-Yang energy distribution with zero volume integral
4.1 Static equilibrium of Euclidean matterspace